# Tracing Multilingual Representations in LLMs with Cross-Layer Transcoders

### v1

@@ 4.2 Mechanisms of Language Decoding
A central question in multilingual models with pivot languages is how models encode language information and integrate it in final layers to predict the correct next token. Across all graphs, language-specific features appear predominantly in early and late layers. In early layers, these features often correspond to single-token content rather than language alone. As in Lindsey et al. (2025b), we find early-layer features indicating output language for specific contexts (e.g., after quotation marks in “The opposite of ‘‘men’’ is ‘‘”), but these features are not consistently present in general cases.
@@
Across all attribution graphs, we find that late-layer language clusters drive decoding. These clusters correspond to high-frequency features that activate for large proportions of tokens within their respective languages. As shown in Appendix C, Figure 13, most high-frequency CLT features are language-specific. Figure 5 shows that these late-layer activations can be linearly read from early-layer features and embedding nodes, indicating that language identity propagates hierarchically through the model.
@@
Finally, activation frequency patterns reveal both language ambiguity and dominance effects. As shown in Figure 6, a small number of features activate on 50–100% of tokens within each language, while inactive tokens occur mostly at sequence beginnings where language identity is ambiguous (especially between French, English, and German). Dominant languages exhibit fewer high-frequency features than others across both GPT-2 and LLaMA models, suggesting that English functions as a default output language requiring fewer explicit high-frequency features Lindsey et al. (2025b).
@@ 4.3 Failure Mechanisms in Non-English Languages
Our analysis in Section 4.1.3 showed that multilingual LLMs form shared representations in their middle layers, suggesting that equivalent meanings are processed through common internal mechanisms across languages. Yet, English-dominant models consistently perform better in English. If the model relies on shared circuits, why does this gap persist?
To answer this, we study two tasks probing distinct reasoning types:
  1. 1.
  2. 2.
Figure 7: Circuit structures across language mixtures. Green: 90% English mixture, orange: 50% mixture. Key clusters missing in Arabic at 90% emerge by 50%, indicating mixture-dependent circuit formation. Additional examples appear in Appendix G.
4.3.1 Antonym Task
Circuits driving antonym prediction are consistent across languages and mixtures, but their completeness depends on mixture balance. In Arabic under the 90% English mixture, the Men&Women cluster, present in most other languages, is missing but appears from the 50% mixture onward (Figure 7). Intervention experiments (Figure 8) confirm its causal role: adding this missing cluster restores correct Arabic predictions (Appendix H).
We also find that the 20% mixture performs best for Latin-script languages (English, French, Arabic), as they share more transferable features than others (Appendix K).
4.3.2 Category Completion Task
We analyze the Category Completion task using the 20% mixture to understand performance in the balanced setting. Arabic performs poorly due to tokenization fragmentation, which forces early layers to reconstruct words rather than activate answer-related clusters. Intervention experiments (Appendix I) show that scaling these clusters restores correct predictions, emphasizing the importance of activation strength.
Comparatively, English consistently triggers stronger responses in task-specific circuits (Men&Women for Antonym, Say Sports for Category Completion). This difference narrows under balanced mixtures but remains linked to sub-tokenization: languages with fragmented tokens show weaker edge strengths from embeddings to target clusters, partially explaining their reduced accuracy (Appendix J).
5 Implications
CLTs make it possible to trace how multilingual models process information step by step, from early encoding to shared representations and finally to language-specific decoding. By exposing feature interactions across layers, they reveal which parts of this pathway behave similarly across languages and where they start to diverge. This provides a concrete way to investigate language disparities: instead of relying only on task performance, we can directly observe which circuits are weaker, missing, or more fragmented for certain languages. More broadly, the ability to follow these internal mappings offers a structured foundation for future work on multilingual models, helping researchers identify which aspects of multilingual processing are stable and which are responsible for cross-language differences.
@@ 6 Conclusion
Model Scale and Architecture. Our primary analysis focuses on relatively small models (68M and 177M parameters) trained from scratch. While we include LLaMA-1B for validation and observe consistent patterns, our findings may not fully generalize to production-scale models (10B+ parameters) or models trained with different architectural choices. Larger models may develop additional mechanisms or more nuanced multilingual representations that our current analysis does not capture.
Performance Disparity Analysis. Our investigation into why English outperforms other languages (Section 4.3) is based on a limited set of case studies across specific tasks (antonym and category completion). While these examples provide mechanistic insights into failure modes, they represent only a small sample of possible multilingual interactions. The findings, particularly regarding circuit activation strength and sub-tokenization effects, may not generalize to all task types or capture the full complexity of performance disparities across languages. A more comprehensive analysis would require systematic evaluation across diverse linguistic phenomena and larger example sets.
@@
Ethical Considerations
Our analysis reveals systematic performance disparities across languages despite balanced training data. Arabic suffers from heavy sub-tokenization, forcing early layers to reconstruct fragmented words rather than process semantics, a structural disadvantage affecting billions of non-Latin script users. While we release our models and CLTs for reproducibility, training required thousands of GPU-hours, raising environmental concerns. Our intervention experiments demonstrate direct manipulation of model outputs across languages. Though intended for improving fairness, these techniques could potentially be misused. Our study covers only five languages and small models (68M, 177M and 1B parameters), limiting generalizability to the world’s  7,000 languages and production-scale systems. We hope these mechanistic insights contribute to more equitable multilingual AI development.
@@ Acknowledgment
We thank Michael Hanna for insightful discussions on the added value of Cross-Layer Transcoders compared to Single-Layer Transcoders, as well as on the training challenges and implementation details of CLTs. We are grateful to Anson Lei for his assistance with the trained GPT-2 model generation issue and for sharing tips on prompt formatting. We also thank Francesco Ortu for his valuable feedback, which helped improve the clarity and flow of the paper.
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by Schmidt Sciences SAFE-AI Grant; by the Frontier Model Forum and AI Safety Fund; by Open Philanthropy; by the Cooperative AI Foundation; and by the Survival and Flourishing Fund. Resources used in preparing this research project were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. F. D. acknowledges support through a fellowship from the Hector Fellow Academy.
@@ A.2.2 Training Scale and Compute Optimization
Following Chinchilla scaling laws Hoffmann et al. (2022), we compute optimal training data sizes:
@@ A.2.3 Training Implementation
We use the nanoGPT codebase (Karpathy, 2022) for reproducible transformer training.
@@
Parameter Value
Optimizer AdamW
Learning Rate 6e-4
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
Weight Decay 0.1
Gradient Clipping 1.0
Warmup Steps 2000
LR Schedule Cosine linear warmup
Batch Size 64
Gradient Accumulation 8 steps
@@
Table 4: Training hyperparameters.
@@
  • Hardware: 4×H100 80GB GPUs per model
  • Distributed Training: Data parallel with gradient synchronization
  • Mixed Precision: FP16 training with automatic loss scaling
  • Checkpointing: Every 5000 steps with automatic resumption
  • Monitoring: Real-time loss and gradient norm tracking
@@ A.2.4 Training Dynamics
Figure 2 shows training and validation loss across English data proportions. Higher English proportions reduce English validation loss but can slightly degrade performance on other languages.
A.2.5 Training Procedure and Model Matrix
Each architecture is trained on all data mixtures, producing a systematic model matrix:
Architecture 20% 50% 70% 90%
GPT-2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TinyStories $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Total Models: 8
Table 5: Complete model training matrix.
For comparative purposes, we also include a pre-trained LLaMA-1B model alongside our custom-trained models.
@@ B.1 Training Configuration
We report in Table 6, the hyper-parameters of the CLT training. We mostly follow Anthropic guidelines Ameisen et al. (2025). This setup is similar across Tinystories, GPT-2, and Llama CLTs.
Table 6: Training Configuration
Parameter Value
@@
Input dimension ($d_{\text{in}}$) 768
Latent dimension ($d_{\text{latent}}$) 24,576
Expansion factor 32
Context size 16
@@
Learning rate $2\times 10^{-4}$
Adam $\beta_{1}$ / $\beta_{2}$ 0.9 / 0.999
Batch size (tokens) 1,024
LR warm-up steps 1,000
@@
Bandwidth 1.0
Initial threshold 0.03
@@
Seed 42
Precision float32 (mixed)
@@ B.2 Performance
We report the amount of dead features and the performance of the different CLTs in Figures  9  10  11.
@@
Appendix C Language Features

### v2

@@ 4.2 Mechanisms of Language Decoding
A central question in multilingual models with pivot languages is how models encode language information and integrate it in final layers to predict the correct next token. Across all graphs, language-specific features appear predominantly in early and late layers. In early layers, these features often correspond to single-token content rather than language alone. As in (Lindsey et al., 2025b), we find early-layer features indicating output language for specific contexts (e.g., after quotation marks in “The opposite of ‘‘men’’ is ‘‘”), but these features are not consistently present in general cases.
@@
Across all attribution graphs, we find that late-layer language clusters drive decoding. These clusters correspond to high-frequency features that activate for large proportions of tokens within their respective languages. As shown in Appendix D, Figure 18, most high-frequency CLT features are language-specific. Figure 4 shows that these late-layer activations can be linearly read from early-layer features and embedding nodes, indicating that language identity propagates hierarchically through the model.
@@
Finally, activation frequency patterns reveal both language ambiguity and dominance effects. As shown in Figure 5, a small number of features activate on 50–100% of tokens within each language, while inactive tokens occur mostly at sequence beginnings where language identity is ambiguous (especially between French, English, and German). Dominant languages exhibit fewer high-frequency features than others across both GPT-2 and LLaMA models, suggesting that English functions as a default output language requiring fewer explicit high-frequency features (Lindsey et al., 2025b).
Language-specific features exhibit significant overlap with English features that increases with English training proportion (77% to 84%), while non-English languages maintain stronger mutual feature overlap ($89\%$) that progressively converges toward English at higher proportions. This pattern indicates that the shared multilingual space emergence is partly driven by training data composition, with English acting as a convergence point (Appendix D).
@@ 4.3 Failure Mechanisms in Non-English Languages
As shown in Section 4.1.3, multilingual LLMs develop shared representations in their middle layers, indicating that equivalent meanings are processed via common internal mechanisms across languages. Despite this, models consistently perform better on English, and even models trained on uniformly distributed multilingual data show a persistent English bias. This raises a key question: if the model relies on shared circuits, why does this performance gap persist?
To address this, we analyze failure and repair mechanisms in non-English languages using a model-diffing approach. Intuitively, the features a model adjusts to perform better in a language reveal what it struggles with most. Let $\mathcal{M}_{0}$ denote the base multilingual model, and $\mathcal{M}_{\ell}$ a version finetuned on language $\ell$. We finetune the original CLT (trained on activations from $\mathcal{M}_{0}$) on the activations of $\mathcal{M}_{\ell}$ to identify the features the model modifies to improve performance. We denote the corresponding feature dictionaries as $\mathcal{D}_{0}$ and $\mathcal{D}_{\ell}$.
We present the findings from our model-diffing experiments on low-performing languages $\ell\in\{\texttt{Arabic, Chinese, German}\}$. Our goal is to understand which features and layers $\mathcal{M}_{\ell}$ adjusts relative to the base model $\mathcal{M}_{0}$, and how these changes propagate to token-level representations.
4.3.1 Models Repair in Later Layers
For each language $\ell$, we compute the cosine similarity between the feature dictionaries $\mathcal{D}_{0}$ and $\mathcal{D}_{\ell}$ across layers.
Figure 6 shows that the largest changes occur in the later layers. This finding aligns with our previous observation that language-specific features are predominantly located in these layers. It is intuitive that the model focuses on refining these representations: by strengthening their connection to the multilingual shared space in the middle layers, $\mathcal{M}_{\ell}$ can more effectively translate and map language-specific inputs to shared high-level representations, ultimately improving performance.
Figure 6: Cosine similarity between $\mathcal{D}_{0}$ and $\mathcal{D}_{\ell}$ across layers for Arabic (left). German (middle) and Chinese (right). The largest representation changes occur in the late layers, indicating that the model focuses its most transformative adaptations there.
4.3.2 Repair Through Multilingual Feature Changes
To understand how finetuning affects multilingual representations, we split features in each layer into two buckets based on cosine similarity: high-change features ($\text{cos-sim}(\mathcal{D}_{0},\mathcal{D}_{\ell})<0.5$) and low-change features ($\text{cos-sim}(\mathcal{D}_{0},\mathcal{D}_{\ell})\geq 0.5$).
Figure 7 shows the distribution of high-change features across layers. Across all languages, high-change features are concentrated in late layers, consistent with our previous analysis. Interestingly, we observe additional patterns in the early and late layers:
  • Layer 0 (embedding layer): Features remain largely unchanged, indicating that the model preserves the base token embeddings.
  • Layer 1: Changes primarily eliminate reductions for Chinese, while English and French dominate the remaining high-change features. This suggests early-layer adjustments help align language-specific signals with shared representations.
  • Layer 11 (late layer): Most transformative changes occur here. High-change features (bucket 1) consistently increase probabilities for the target language across all languages, indicating that the model performs subtle adjustments—minor rotations or rescaling—to align language-specific features with the correct output while maintaining stable, well-behaved representations.
These adjustments naturally push features from other languages into bucket 2 (low similarity), balancing the representation space. Notably, we also observe cross-language effects: finetuning German increases Arabic features in bucket 1, and Chinese exhibits similar interactions with French and Arabic. This hints at overlapping subspaces between languages, where modifications in one language can partially transfer to others. More results and experimental setup are shown in Appendix J.1.
Figure 7: Language distribution comparison across layers for German. Middle layers show high entropy and low language distribution deltas and are therefore omitted from the figure. Results for Arabic and Chinese are presented in Appendix J.1
4.3.3 Causal Impact of High-Change Features
To assess the functional role of the high-change features identified in $\mathcal{D}_{\ell}$, we perform a causal analysis using LogitLens. For a given high-change feature $f\in\mathcal{D}_{\ell}$ and $\mathcal{D}_{0}$, we compute its top aligned tokens via the unembedding matrix:
$\mathbf{t}^{*}(\mathbf{f})=\arg\max_{i}\big[W_{U}W_{\text{dec}}f\big]_{i},$
where $W_{\text{dec}}$ is the decoder projection and $W_{U}$ is the unembedding matrix. This operation identifies the tokens most strongly influenced by $\mathbf{f}$.
We find that, after finetuning, these features contribute to increased token assembly in early layers (see Figure 43). This suggests that late-layer adaptations in $\mathcal{D}_{\ell}$ propagate backward, enabling $\mathcal{M}_{\ell}$ to construct more abstract, compositional representations. Furthermore, the norms of high-change features in $\mathcal{D}_{\ell}$ are larger than in $\mathcal{D}_{0}$, reflecting stronger and more confident activations.
For further details on the features analyzed and the results of the Logit Lens experiment, see Appendix J.2.
4.3.4 Circuit-Level Illustration
To illustrate our findings, we present over 40 representative circuit examples from Arabic, Chinese, and German (full details in Appendix G.2). These examples show how interventions that activate clusters of high-change features can improve the model’s retrieval of the correct answer. We further observe that sub-tokenization weakens edges to the most important clusters, and that the tokenizer—despite being trained on equal amounts of data per language—exhibits a bias toward English. This bias provides a mechanistic explanation for the performance gaps observed across languages.
@@ 5 Conclusion
Model Scale and Architecture. Our primary analysis focuses on relatively small models (68M and 177M parameters) trained from scratch. While we include LLaMA-3.2-1B for validation and observe consistent patterns, our findings may not fully generalize to production-scale models (10B+ parameters) or models trained with different architectural choices. Larger models may develop additional mechanisms or more nuanced multilingual representations that our current analysis does not capture.
@@
Impact Statement
CLTs make it possible to trace multilingual processing step by step—from early encoding through shared representations to language-specific decoding—revealing where mechanisms are shared across languages and where they diverge. This enables a concrete, mechanistic view of language disparities: rather than relying solely on task-level performance, we can directly identify circuits that are weaker, missing, or more fragmented for particular languages. Using this lens, we show that even under balanced training data, multilingual models exhibit systematic gaps; for example, Arabic is heavily sub-tokenized, requiring early layers to reconstruct fragmented words. We release our models and CLTs to support further research, while noting that training required thousands of GPU-hours and that intervention capabilities raise fairness and misuse considerations. Although our experiments focus on five languages and relatively small models, the approach provides a structured foundation for analyzing multilingual processing and developing more equitable multilingual AI systems.
@@ Acknowledgment
We thank Michael Hanna for insightful discussions on the added value of Cross-Layer Transcoders compared to Single-Layer Transcoders, as well as on the training challenges and implementation details of CLTs. We are grateful to Anson Lei for his assistance with the trained GPT-2 model generation issue and for sharing tips on prompt formatting. We also thank Francesco Ortu for his valuable feedback, which helped improve the clarity and flow of the paper and Antia Garcia for helping improve the design of the main figure.
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by Schmidt Sciences SAFE-AI Grant; by the Frontier Model Forum and AI Safety Fund; by Open Philanthropy; by the Cooperative AI Foundation; and by the Survival and Flourishing Fund. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. F.D. acknowledges support through a fellowship from the Hector Fellow Academy. We gratefully acknowledge EleutherAI for providing compute resources that enabled the Cross-Layer Transcoder finetuning in our model-diffing experiments.
Supplementary Materials
PART I: Additional Methodological Details .A
A Multilingual Model Training A
A.1  Tokenization Strategy A.1
A.2  Model Architectures and Training A.2
B CLT Training B
B.1  Training Configuration B.1
B.2  Performance B.2
C Additional Validation Loss Curves Under Data Imbalance C
C.1  GPT2 Models C.1
C.2  TinyStories Models C.2
D Language Feature Analysis D
D.1  Methodology and Definitions D.1
D.2  Feature Overlap Results D.2
PART II: Complementary Analyses .E
E Multilingual Score E
F Shared Multilingual Space Independence from English Dominance F
PART III: Case Studies .G
G Case Studies: Multilingual Circuits G
G.1  Multilingual Circuits: U-shaped Entropy G.1
G.2  Multilingual Circuits: Failure Modes G.2
G.3  Intervention Validation G.3
H Extended Task Analysis: Translation and Cultural Prompts H
H.1  Translation Prompts H.1
H.2  Cultural Context Prompts H.2
I Why English Performs Better: A Case Study I
I.1  Cluster Activation Analysis I.1
I.2  Sub-tokenization Effects I.2
PART IV: Model Diffing and Tokenizer Analysis .J
M Model Diffing: Additional Material J
M.1  Multilinguality of Repair Features J.1
M.2  Token Assembly in Early Layers J.2
N Tokenizer Analysis: BPE Constraints on Multilingual Processing K
N.1  Vocabulary Allocation: Cannibalization Testing K.1
N.2  Morphological Coherence K.2
N.3  Mechanistic Implications K.3
@@ A.2.2 Training Scale and Compute Optimization
Following Chinchilla scaling laws (Hoffmann et al., 2022), we compute optimal training data sizes in Table 3.
@@ A.2.3 Training Implementation
We use the nanoGPT codebase (Karpathy, 2022) for reproducible transformer training and use the hyperparameters presented in Table 4.
@@
Parameter Value Parameter Value
Optimizer AdamW Learning Rate 6e-4
$\beta_{1}$ 0.9 $\beta_{2}$ 0.95
Weight Decay 0.1 Gradient Clipping 1.0
Warmup Steps 2000 LR Schedule Cosine linear warmup
Batch Size 64 Gradient Accumulation 8 steps
@@
Table 4: Training hyperparameters (horizontal layout for one-column appendix).
@@
We train each model on 4×H100 80GB GPUs. Training is distributed using data parallelism with gradient synchronization, and we use mixed precision (FP16) with automatic loss scaling.
@@ A.2.4 Training Dynamics
Figures 11, 12 and 13 shows training and validation loss across English data proportions. Higher English proportions reduce English validation loss but can slightly degrade performance on other languages.
A.2.5 Training Procedure and Models Trained
Each architecture is trained on all data mixtures: $20\%,50\%,70\%$ and $90\%$. Furthermore, we train another model on non-English data ($25\%$ for each of Arabic, Chinese, German and French).
For comparative purposes, we also include a pre-trained LLaMA-3.2-1B (Meta AI, 2026) model alongside our custom-trained models.
@@ B.1 Training Configuration
We report in Table 5, the hyper-parameters of the CLT training. We mostly follow Anthropic guidelines (Ameisen et al., 2025). This setup is similar across Tinystories, GPT-2, and Llama CLTs.
Parameter Value Parameter Value
@@
Input dimension ($d_{\text{in}}$) 768 Latent dimension ($d_{\text{latent}}$) 24,576
Expansion factor 32 Context size 16
@@
Learning rate $2\times 10^{-4}$ Adam $\beta_{1}$ / $\beta_{2}$ 0.9 / 0.999
Batch size (tokens) 1,024 LR warm-up steps 1,000
@@
Bandwidth 1.0 Initial threshold 0.03
@@
Seed 42 Precision float32 (mixed)
Table 5: Training Configuration (horizontal layout for one-column appendix).
@@
For all CLTs, we stop training when the L0 sparsity coefficient goes below 10 and Explained Variance is beyond $70\%$. At this point, we save the full model activations and then load them in random chunks from the dataset for training. The CLTs contain roughly $78\times 768\times 32\times 768$ parameters, though for the LLaMA CLT we use fewer features (expansion factor = 4) due to GPU memory constraints. Replacement scores are consistently in the range 0.75–0.8, which aligns with expected reconstruction performance. For circuit extraction, attribution graphs are pruned to retain 80% of cumulative node scores and 85% of cumulative edge scores using the circuit-tracing library. Training is conducted over approximately 200M tokens and follows standard CLT guidelines (Ameisen et al., 2025).
@@ B.2 Performance
We report the amount of dead features and the performance of the different models (TinyStories, GPT2, Llama-3.2-1B) CLTs in Figures  8,  9 and  10.
@@
Appendix C Additional Validation Loss Curves Under Data Imbalance
C.1 GPT2-Models
In Section 4.1.1, we reported that the validation loss curves for the four non-English languages (Chinese, Arabic, German, French) show competitive performance. Here we show the curves.
We also provide additional results for Chinese . Figure 12 shows the validation loss curves for Chinese. The results mirror the patterns observed for other languages: despite extreme imbalance in the training data, the model maintains competitive prediction performance in Chinese, indicating that the robustness of multilingual circuits extends to typologically distinct languages.
C.2 TinyStories Models
Appendix D Language Feature Analysis
We analyze whether high-frequency features exhibit language specificity and quantify the extent to which language-specific features are shared across languages. This analysis uses precomputed activation statistics from CLT training with varying English proportions in the training data: 20%, 50%, 70%, and 90% English.
D.1 Methodology and Definitions
For each feature $f$ at layer $\ell$, we compute two statistics from the training data. First, the frequency $\text{freq}(f,\ell)$ represents the proportion of tokens for which the feature activates. Second, the language distribution $P(\text{lang}\mid f,\ell)$ captures the probability distribution over languages conditioned on feature activation. Specifically,
$P(\text{lang}\mid f,\ell)=\frac{N_{\text{seq}}(f,\ell,\text{lang})}{N_{\text{seq}}(f,\ell,\text{total})},$
where $N_{\text{seq}}(f,\ell,\text{lang})$ denotes the number of sequences in language lang in which feature $f$ activates at least once.
To label a feature as language-specific, the feature must firstly be high-frequency, with $\text{freq}(f,\ell)\geq 0.03$. Second, the feature must be statistically associated with a given language, requiring $P(\text{lang}\mid f,\ell)\geq 0.5$, i.e., at least 50% of the sequences on which it activates are from that language. For example, a feature $f$ at layer $\ell$ is labeled French-specific if both $\text{freq}(f,\ell)\geq 0.03$ and $P(\text{French}\mid f,\ell)\geq 0.5$.
For each language pair, we compute feature overlap via set intersection. Let
$F_{\text{lang\_a}}=\{(\ell,f):\text{freq}(f,\ell)\geq 0.03\land P(\text{lang\_a}\mid f,\ell)\geq 0.5\}$
denote the set of Language-A-specific features across all layers, with an analogous definition for $F_{\text{lang\_b}}$. We define overlap as
$\frac{|F_{\text{lang\_a}}\cap F_{\text{lang\_b}}|}{|F_{\text{lang\_a}}|}\times 100\%,$
D.2 Feature Overlap Results
Figure 14 shows the percentage of each non-English language’s language-specific features that overlap with English-specific features across the four training checkpoints. We observe a consistent increase in overlap as the proportion of English training data increases, rising from approximately 77% at 20% English to 84% at 90% English across all non-English languages. This trend indicates that increasing English exposure during training drives convergence toward an English-dominant feature space.
Figure 14: Overlap between non-English language–specific features and English-specific features across training checkpoints.
Figure 15 presents pairwise feature overlap among all non-English language pairs. We find that non-English languages overlap more strongly with each other (averaging 89% at 20% English) than they do with English (averaging 77% at 20% English), yielding a gap of approximately 12 percentage points. However, this gap progressively closes as English training proportion increases, reaching near parity at 90% English, indicating forced convergence.
Figure 15: Pairwise overlap of language-specific features among non-English languages.
Figure 16 shows the full pairwise overlap matrix for the 90% English checkpoint as a heatmap. The visualization reveals notable asymmetries in pairwise relationships. For example, Chinese and Arabic maintain 86.6% overlap, substantially higher than the 78.1% overlap observed between French and German. All pairwise overlaps exceed 78%, confirming the existence of a broadly shared feature space. English exhibits relatively symmetric overlap with all other languages (83–84%), consistent with its role as a convergence point under high training proportions.
Figure 16: Complete pairwise feature overlap matrix at 90% English training.